# suppSupplement

### v1

@@ Abstract
In optics and photonics, a small number of building blocks, like resonators, waveguides, arbitrary couplings, and parametric interactions, allow the design of a broad variety of devices and functionalities, distinguished by their scattering properties. These include transducers, amplifiers, and nonreciprocal devices, like isolators or circulators. Usually, the design of such a system is handcrafted by an experienced scientist in a time-consuming process where it remains uncertain whether the simplest possibility has indeed been found. In our work, we develop a discovery algorithm that automates this challenge. By optimizing the continuous and discrete system properties our automated search identifies the minimal resources required to realize the requested scattering behavior. In the spirit of artificial scientific discovery, it produces a complete list of interpretable solutions and leads to generalizable insights, as we illustrate in several examples. This now opens the door to rapid design in areas like photonic and microwave architectures or optomechanics.
April 30, 2024
@@
Waves represent one of the most basic physical phenomena. This explains the importance of applications enabled by wave transport in suitably designed structures, ranging from integrated photonics to microwave circuits. A large class of on-chip devices, from isolators and circulators for non-reciprocal transport [1, 2, 3, 4] to hybrid frequency converters [5, 6] and new classes of amplifiers [7, 8], can be engineered by combining the right building blocks in a well-chosen scattering topology. So far, each new design in this domain has been proposed based on human ingenuity and experience. It would be desirable to speed up the pace of this laborious exploration process in the high-dimensional space of possible setups and obtain a complete overview of all conceptually distinct options for any given desired functionality. Here, we introduce an automated discovery algorithm that achieves these goals and also helps to uncover new insights.
Tackling challenges of this kind is the domain of artificial scientific discovery [9]. This rapidly evolving field has the ambitious aim to automate all aspects of the scientific process using tools of machine learning and artificial intelligence, with an emphasis on interpretability and discovery of new conceptual insights. Pioneering examples are the identification of organic molecules from mass spectrometry by the computer program Dendral [10] and the generation and experimental testing of hypotheses in biochemistry by a robot scientist [11]. Important challenges in this field include automatic extraction of symbolic expressions [12, 13], discovery of collective coordinates [14, 15], automated planning and execution of experiments [16, 17], and generating new experimental setups. The latter task has been successfully demonstrated for the preparation of multi-photon states in quantum optics [18, 19], for the optimization of superconducting circuits [20], gravitational wave detectors [21], and signal processing using optical components [22].
Automated discovery relies crucially on casting the complex search space into a unifying language. For the domain envisaged here, we identified coupled-mode theory as providing a sweet spot on the continuum between more microscopic hardware-specific descriptions (such as engineered refractive-index distributions [23, 24]) and descriptions based on macroscopic building blocks on the level of entire functional elements (like amplifiers). The latter require a larger number of building blocks and need more human design input. For example [22] focuses on signal processing setups that can be decomposed into directed networks of specific technical components. In contrast, coupled-mode designs are transferable between different platforms and can always be translated back into concrete physical structures. Of equal importance is the formalization of the goal. Here, we chose a description in terms of the scattering matrix, specifying the desired functionality via the externally visible characteristics.
Our discovery algorithm is based on an efficient search process that exploits the recursive modular structure of the problem to prune the exponentially large space of possibilities and provide a complete list of setups fulfilling specified constraints. It automatically ends up with interpretable and generalizable results both by using discrete building blocks and by suggesting only dimensionless parameters of universal, implementation-independent meaning. The final outcome is a menu of irreducible setups that allows to select optimality criteria even afterwards. Moreover, we introduce a fruitful general extension of the toolbox available in artificial discovery, observing that conceptual insights in design are often based on thinking about idealized structures realizing asymptotic limits of parameters. We describe a modification of our algorithm that exploits this meta-principle.
We illustrate the benefits of this approach for a variety of target functionalities, including couplers, circulators and directional amplifiers, where it gives rise to new discoveries and generalizable insights.
@@
Figure 1: Graph representation of coupled mode systems. (a) Physical modes correspond to graph nodes, couplings to edges. Port modes (orange) are connected to waveguides which relay input and output signals. (b) Discrete building blocks available to the discovery algorithm. (c) Illustration of hardware platforms which are efficiently described by coupled mode theory. This includes phontonic crystals (left), optomechanical setups (top), and microwave setups (right).
We consider a network of coupled bosonic modes, which can be implemented as an optical, electric, mechanical, or hybrid setup, comprising different types of modes. This device is probed with multiple input signals. We assume that the interaction between the modes can be engineered within some constraints, set by the envisioned experimental implementation. Our method takes into account these constraints but also that certain types of interaction might be more difficult to realize than others, which is factored in the definition of the setup complexity. The choices for the mode and interaction types produce a large discrete search space growing exponentially with the number of modes required for a target response.
@@
The linear response of a multi-mode circuit to signals of a given frequency is encoded in its scattering matrix $S$. The element $S_{ji}$ is the transmission coefficient of a signal from the input port $i$ to the output port $j$. In our work, we aim to automatically discover the simplest setups that fulfill a certain target response, which is encoded in a target scattering matrix $S_{\mathrm{target}}$.
@@
The precise characterization of a specific physical implementation of an ideal target device requires a carefully crafted system-specific model. In contrast, the ideal behavior is typically captured by a simple and general high-level description. Like a human scientist would do, we choose our artificial scientist to adopt such high-level description. This approach allows us to discover ideal and transferable solutions.
@@
At the highest level of description, any multi-mode circuit can be represented as a graph with colored edges [25] (see Fig. 1(a,b)), regardless of the underlying hardware platform (see examples in Fig. 1(c)). Nodes correspond to modes and edges correspond to physical interactions. “Port modes” (orange) are connected to waveguides transmitting the input and output signals of interest. We note that the carrier frequencies $\omega_{L,j}$ of the input fields should be close to the respective mode resonances $\omega_{j}$ and can be different for different port modes, e.g. an optical and a mechanical mode will have very different frequencies. We focus on a fixed set of frequencies, but our approach could be extended to also consider aspects like optimizing the bandwidth [26].
@@
The port modes are often insufficient to implement a given scattering matrix. Therefore, we introduce a variable number of auxiliary modes (grey) interacting with the port modes via the available interactions.
A linear multi-mode scattering setup can contain several interactions: passive coupling (black), nonreciprocal coupling (green) acquiring a direction-dependent phase, and squeezing (blue). It is useful to distinguish between these different types of quadratic interactions because in relevant experimental scenarios, they require qualitatively different implementations. Passive coupling can be realised, e.g. by evanescent coupling between defects modes in a photonic crystal. In contrast, the other interactions are typically generated by pumping a weak higher-order coupling, e.g. using a Kerr medium or optomechanics, or, for nonreciprocal coupling with a directional phase, by coupling modes that are not invariant under the time-reversal symmetry, e.g. whispering gallery modes [27]. In any case, all of these interactions can be written in the form $g_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}+h.c.$ (with $g_{ij}$ real-valued for passive coupling) or $\nu_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}+h.c.$, respectively (using a rotating wave approximation; see Supplemental Material). Here, $h.c.$ denotes the Hermitian conjugate, and $\hat{a}^{\dagger}_{j}$ excites mode $j$. The coupling rates $\nu_{ij}$ can be viewed as elements of a symmetric matrix, $g_{ij}$ as elements of a Hermitian matrix. In our description, any pair of off-diagonal non-zero elements, $\nu_{ij}\neq 0$ or $g_{ij}\neq 0$, corresponds to a graph edge.
@@
Figure 2: Automated Discovery of Modular Scattering Setups. (a) Optimisation scheme. The discrete optimization sorts graphs into the libraries of valid (green book) and invalid graphs (red book). In every round, a new graph is suggested and, depending on the outcome of the continuous optimisation, sorted with all its extensions/subgraphs into the library of valid/invalid graphs. The continuous optimization minimizes the loss function to realize a target scattering behaviour (left block), here an isolator. The candidate graph (right block) defines the number of auxiliary modes and an allowed set of non-zero couplings rates and phases in the dimensionless Hamiltonian $H$. (b) Search space (schematic), starting with a fully connected graph and successively pruning edges, resulting in multiple irreducible solutions. (c) Comparison of the number of tested graphs for the directional coupler (see Fig. 3(b)), which uses five modes. The total number of possible graphs (black) for a certain number of free variables in $H$ is compared to the number of graphs which had actually to be tested by our exhaustive search. Our scheme starts with maximally connected graphs and successively goes to more restricted graphs. By testing only around 13,000 graphs we are able to characterize all 1.9 million graphs in this example.
@@
To identify setups that implement the target scattering matrix, we have to explore the discrete search space of all possible graphs and, given a graph, find an appropriate set of values for the underlying continuous parameters, i.e., the non-zero couplings, phases, and loss rates. To this end, we perform a two-step procedure: A discrete optimization routine suggests new graphs and an embedded continuous optimization algorithm looks for appropriate values of the parameters (see Fig. 2(a)). If the continuous optimization is successful, we label a graph as valid.
@@
First, we discuss the discrete optimization, see Fig. 2(a). We aim to identify the simplest valid graphs as they represent the simplest experimental implementations of the target response. Often there exist a multitude of valid graphs that differ in the number and the complexity of the required couplings. Our algorithm identifies “irreducible” graphs that can not be further simplified by setting any coupling rate or phase to zero. To identify all such graphs we perform a complete enumeration.
@@
Due to the large discrete search space, a brute-force approach testing all graphs is infeasible in many cases of interest. We solve this problem by leveraging the knowledge gained from previously tested graphs, drastically reducing the required computational resources. Note that an alternative approach could be developed based purely on continuous optimization, introducing terms in a loss function that reward sparsity (zero couplings) as in [28, 29, 30] – however, this would not enable us to obtain a complete set of solutions.
@@
We take advantage of the fact that removing an edge from a graph is equivalent to setting a coupling rate to zero. Thus, if a graph is found to be invalid, that carries over to all graphs which can be generated by removing some edges and/or setting coupling phases to zero (changing an edge color from green to black). Likewise, all extensions of a valid graph are valid (see Fig. 2(b)). This observation allows us to efficiently construct exhaustive libraries of valid and invalid graphs by testing only a small fraction, see Fig. 2(c).
First, we test only fully connected graphs for validity, until we identify the minimum required number of modes. Starting from that parent graph, our algorithm suggests at every iteration a new graph that is not yet included in any of the two libraries and tests for validity. If no solution is discovered by the continuous optimization, the graph and all its subgraphs are added to the library of invalid graphs. Otherwise, the graph and all its extensions are added to the library of valid graphs. If some of the discovered parameters turn out to be zero, the corresponding simpler graph and all its extensions are added instead.
@@
We now turn to the continuous optimization. The scattering matrix $S$ is fully determined by an appropriately defined first-quantized dimensionless Hamiltonian $H$ (see below), which contains the free coupling parameters. We look for setups implementing the desired scattering behavior by minimizing the square deviation loss function with respect to those couplings:
@@
We note that the target scattering matrix $S^{\mathrm{target}}$ can contain further free parameters that are not specified by the desired scattering properties. For example, we might impose that two scattering amplitudes are equal but a suitable value needs to be discovered. In addition, we introduce the free parameters $\gamma^{\rm in/out}_{i}$ to account for the fact that the optical path length of the input and output fields in the waveguide can be easily modified. Finally, we allow for extra constraints $f_{j}$ that can be expressed in terms of the multi-mode circuit parameters or the scattering matrix $S$, e.g. when enforcing minimum added noise. We minimize the loss ${\cal L}$ as a function of the free parameters in $H$, $S_{\mathrm{target}}$, and the phases $\gamma^{\rm in/out}_{i}$. For the minimization, we use the Broyden–Fletcher–Goldfarb–Shanno algorithm [31], an iterative gradient-descent method for solving non-linear optimization problems. A valid solution is obtained whenever the loss reaches zero. The algorithm fails to find a solution if the loss gets stuck in a non-zero minimum. Since the optimization landscape supports local minima, there is a risk of false-negative errors. To reduce this risk, we repeat the continuous optimization $N_{\rm rep}$ times. The hyperparameter $N_{\rm rep}$ is chosen empirically.
A key step of our method is to derive the scattering matrix $S$ in terms of an appropriately rescaled Bogoliubov-de-Gennes Hamiltonian:
$H=\frac{1}{\sqrt{\kappa}}\begin{pmatrix}g&\nu\\ \nu^{*}&g^{*}\\ \end{pmatrix}\frac{1}{\sqrt{\kappa}},$ (2)
see Supplemental Material. Here, $\kappa$ is the diagonal matrix $\kappa={\rm diag}(\kappa_{1},\ldots,\kappa_{N},\kappa_{1},\ldots,\kappa_{N})$ with $\kappa_{i}$ being the out-coupling of mode $i$ into the corresponding output port (for an auxiliary mode it could be an intrinsic loss channel). The rescaling in Eq. (2) is important, since it means the algorithm discovers classes of solutions that depend only on the dimensionless detuning $-H_{ii}=(\omega_{L,i}-\omega_{i})/\kappa_{i}$, the cooperativities ${\cal C}_{ij}=4|H_{ij}|^{2}$, and the gauge invariant phases accumulated on closed loops (synthetic field fluxes), $\Phi_{i,j,l,\ldots,k}=\arg(H_{ij}H_{jl}\ldots H_{ki})$. When translating back to a concrete physical setup, this leaves the freedom to choose arbitrarily the decay rates $\kappa_{i}$ and thus the couplings (although for amplifiers some additional stability constraints might apply, see Supplemental Material). For simplicity, we focus below on the important special case of phase-preserving devices. In this scenario, the scattering matrix is an $N\times N$ matrix as already implicitly assumed in Eq. 1. In addition, we have assumed that each port mode is coupled in reflection to a single waveguide. For further details like the treatment of quantum noise, phase-sensitive amplifiers, coupling in transmission, and application to setups with chiral modes, see the Supplemental Material.

### v2

# Automated Discovery of Coupled Mode Setups

@@ Abstract
In optics and photonics, a small number of building blocks, like resonators, waveguides, arbitrary couplings, and parametric interactions, allow the design of a broad variety of devices and functionalities, distinguished by their scattering properties. These include transducers, amplifiers, and nonreciprocal devices, like isolators or circulators. Usually, the design of such a system is handcrafted by an experienced scientist in a time-consuming process where it remains uncertain whether the simplest possibility has indeed been found. In our work, we develop the discovery algorithm AutoScatter that automates this challenge. By optimizing the continuous and discrete system properties our automated search identifies the minimal resources required to realize the requested scattering behavior. In the spirit of artificial scientific discovery, it produces a complete list of interpretable solutions and leads to generalizable insights, as we illustrate in several examples. This opens the door towards automated discovery of scattering setups for photonics, microwaves and optomechanics, with possible future extensions to periodic structures, sensing, and electronic devices.
February 21, 2025
@@
Waves represent one of the most basic physical phenomena. This explains the importance of applications enabled by wave transport in suitably designed structures, ranging from integrated photonics to microwave circuits. A large class of on-chip devices, from isolators and circulators for non-reciprocal transport [1, 2, 3, 4] to hybrid frequency converters [5, 6] and new classes of amplifiers [7, 8], can be engineered by combining the right building blocks in a well-chosen scattering topology. So far, each new design in this domain has been proposed based on human ingenuity and experience. It would be desirable to speed up the pace of this laborious exploration process in the high-dimensional space of possible setups and obtain a complete overview of all conceptually distinct options for any given desired functionality. Here, we present the automated discovery algorithm AutoScatter [9] that achieves these goals and also helps to uncover new insights.
@@
Figure 1: Graph representation of coupled-mode systems. (a) Physical modes correspond to graph nodes, couplings to edges. Port modes (orange) are connected to waveguides which relay input and output signals. Auxiliary modes (gray) have an intrinsic loss channel whose output signal is not monitored. The edges’ weights are dimensionless coupling rates which are rescaled over the loss rates. (b) Discrete building blocks available to the discovery algorithm. (c) Illustration of some of the many hardware platforms which are efficiently described by coupled-mode theory. These include photonic circuits, e.g., based on photonic crystals (left), optomechanical setups (top), and microwave circuit setups (right).
Tackling challenges of this kind is the domain of artificial scientific discovery [10]. This rapidly evolving field has the ambitious aim to automate all aspects of the scientific process using tools of machine learning and artificial intelligence, with an emphasis on interpretability and discovery of new conceptual insights. Pioneering examples are the identification of organic molecules from mass spectrometry by the computer program Dendral [11] and the generation and experimental testing of hypotheses in biochemistry by a robot scientist [12]. Important challenges in this field include automatic extraction of symbolic expressions [13, 14], discovery of collective coordinates [15, 16], automated planning and execution of experiments [17, 18, 19], and generating new experimental setups. The latter task has been successfully demonstrated for the preparation of entangled states in quantum optics [20, 21], for the optimization of superconducting circuits [22], gravitational wave detectors [23], and signal processing using optical components [24].
Automated discovery relies crucially on casting the complex search space into a unifying language. For the domain envisaged here, we identified coupled-mode theory [25, 26] as providing a sweet spot on the continuum between more microscopic hardware-specific descriptions (such as engineered refractive-index distributions [27, 28, 29, 30]) and descriptions based on macroscopic building blocks on the level of entire functional elements (like amplifiers). The latter require a larger number of building blocks and need more human design choices; for example [24] focuses on signal processing setups that can be decomposed into networks of specific technical components. In contrast, coupled-mode designs are transferable between different platforms and can always be translated back into several different concrete physical structures, ranging from photonic crystals to optomechanical setups and microwave circuits. Of equal importance is the formalization of the goal. Here, we chose a description in terms of the scattering matrix, specifying the desired functionality via the externally visible characteristics.
Our discovery algorithm AutoScatter is based on an efficient search process that exploits the recursive modular structure of the problem to prune the exponentially large space of possibilities and provides a complete list of setups fulfilling specified constraints. It automatically ends up with interpretable and generalizable results both by using discrete building blocks and by suggesting only dimensionless parameters of universal, implementation-independent meaning. The final outcome is a menu of irreducible setups that allows to select optimality criteria even afterwards. Moreover, we introduce a fruitful general extension of the toolbox available in artificial discovery, observing that conceptual insights in design are often based on thinking about idealized structures realizing asymptotic limits of parameters. We describe a modification of our algorithm that exploits this meta-principle.
We illustrate the benefits of AutoScatter for a variety of target functionalities, including couplers, circulators and directional amplifiers, where it produces new discoveries and generalizable insights.
I Coupled-mode theory
@@
We consider a network of coupled bosonic modes, which can be implemented as an optical, electric, mechanical, or hybrid setup, comprising different types of modes. This device is probed with multiple input signals and its linear response is encoded in its scattering matrix $S$. The element $S_{ji}$ is the transmission coefficient of a signal from the input port $i$ to the output port $j$. In our work, we aim to automatically discover the simplest setups that fulfill a certain target response, which is encoded in a target scattering matrix $S^{\mathrm{target}}$.
@@
The precise characterization of a specific physical implementation of an ideal target device requires a carefully crafted system-specific model. In contrast, the ideal behavior is typically captured by a simple and general high-level description. Like a human scientist would do, we choose our artificial scientist to adopt such high-level description. This approach allows us to discover ideal and transferable solutions. We discuss in Section VIII how to taylor our discovered solutions to specific hardware platforms.
@@
At the highest level of description, any multi-mode circuit can be represented as a graph with colored edges [31] (see Fig. 1(a,b)), regardless of the underlying hardware platform. Nodes correspond to modes and edges represent resonant interactions between the modes. “Port modes” (orange) are connected to waveguides transmitting the input and output signals of interest. The carrier frequencies $\omega_{L,j}$ of the input fields should be close to the respective mode resonances $\omega_{j}$ and can be different for different port modes, e.g., an optical and a mechanical mode will have very different frequencies. Here, we focus on the response of each port mode at the respective carrier frequency, which is described by a single scattering matrix encoding the ideal behaviour of the device. AutoScatter could be extended to also consider aspects like optimizing the bandwidth [32] but this would lead to less transferable solutions, see Section VIII, and is therefore beyond the scope of this work.
@@
The port modes are often insufficient to implement a given scattering matrix. Therefore, we permit to add a variable number of auxiliary modes (gray). The auxiliary modes are also damped but we do not monitor the corresponding loss channels. Therefore, these channels are not included in the target scattering matrix.
We use colored edges to distinguish between different types of interactions. The edges’ weights are appropriately rescaled dimensionless coupling constants. We choose the rescaling such that the weights fully determine the scattering matrix and, thus, also incorporate the essential information about the relative strength of the different coherent and dissipative processes. This rescaling is an essential step in our method and merits a separate discussion below. Before that we discuss the coherent interactions between the modes of the multimode device.
The response of state-of-the-art multimode devices can be precisely characterized by linearizing the full nonlinear dissipative dynamics about a steady state. In this setting, the interactions between $N$ high-quality-factor modes are well described by a time-independent bosonic quadratic Hamiltonian of the form
$\frac{1}{2}\sum_{ij=1}^{N}\left(g_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}+\nu_{ij}\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}\right)+\mathrm{H.c.}$ (1)
Here, $\mathrm{H.c.}$ denotes the Hermitian conjugate, and $\hat{a}_{j}$ and $\hat{a}^{\dagger}_{j}$ are the ladder operators for mode $j$. Any off-diagonal non-zero coupling constant, $\nu_{ij}\neq 0$ or $g_{ij}\neq 0$, corresponds to a graph edge. Importantly, there are two distinct types of interactions: Squeezing interactions with coupling constants $\nu_{ij}$ (corresponding to blue edges in our graph representation) create pairs of excitations in the modes $i$ and $j$, while beamsplitter interactions with coupling constants $g_{ij}$ exchange excitations between the modes $i$ and $j$. In our optimisation, we also distinguish between complex-valued beamsplitter couplings (green) with $g_{ij}\!\in\!\mathbb{C}$ which involve a complex phase depending on the direction of the exchange and real-valued beamsplitter couplings (black) with $g_{ij}\!\in\!\mathbb{R}$ which are direction independent. It is useful to distinguish between the two different types of beamsplitter interactions in view of identifying couplings that can be implemented passively in many common experimental scenarios, see Section VIII for more details.
An important aspect of the Hamiltonian in Eq. 1 is that it is defined in a rotating frame. By applying a so-called “Rotating-Wave-Approximation” (RWA), we drop the fast oscillating terms describing non-resonant interactions (see Appendix D for more details). In this rotating frame, each mode $j$ rotates at its corresponding carrier frequency $\omega_{L,j}$ (for an auxiliary mode $j$, $\omega_{L,j}$ is the vibration frequency induced by a signal entering the device at any port $l$ with the respective carrier frequency $\omega_{L,l}$). Thus, $-g_{jj}$ is to be interpreted as the detuning $\Delta_{j}$ of the carrier frequency $\omega_{L,j}$ from the mode resonance $\omega_{j}$, $\Delta_{j}=\omega_{L,j}-\omega_{j}$. In our graph representation, any non-zero detuning is visualized as self-loop, see Fig. 1(a). Importantly, the RWA Hamiltonian already captures the ideal behavior of the devices even though it does not directly depend on the resonance frequencies $\omega_{j}$. This is an appealing feature that makes any insights gained using this approximation highly transferable across various platforms. For this reason, it is widely adopted for hand-crafted design of multi-mode circuits [2, 3, 4, 5, 6, 8, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. A more refined description including non-resonant interactions is often used to quantify how a specific implementation deviates from the ideal target behavior. This deviation is typically small for high-quality factor modes ubiquitous in modern quantum optics platforms. Moreover, it is device-specific, thus, incorporating it in our approach would lead to less transferable results.
As anticipated above, a key novelty of our graph representation is our rescaling of the weights. Previous works adopting graph representations of multimode circuits [31, 32] had introduced weights by expressing the coupling constants and other frequency parameters in units of a single base frequency. This is also the most common approach in other numerical and analytical studies of the scattering matrix. In this approach, the number of free parameters/graph weights is reduced by one without losing generality. Instead, we achieve a much larger reduction of the free parameters by incorporating in the graph edges $N$ decay rates (one for each mode). We first express the Hamiltonian in Eq. 1 in matrix form, i.e., in first quantization as a so-called Bogoliubov-de-Gennes (BdG) Hamiltonian. Then, we rescale the BdG Hamiltonian in the following way (see Appendix C):
$H=\frac{1}{\sqrt{\kappa}}\begin{pmatrix}g&\nu\\ \nu^{*}&g^{*}\\ \end{pmatrix}\frac{1}{\sqrt{\kappa}}.$ (2)
Here, $\kappa$ is the diagonal matrix $\kappa={\rm diag}(\kappa_{1},\ldots,\kappa_{N},\kappa_{1},\ldots,\kappa_{N})$, and $\kappa_{i}$ is the decay rate of mode $i$ (out-coupling rate for port modes, intrinsic loss rate for auxiliary modes). Our graph weights are then the entries of this dimensionless BdG Hamiltonian. Importantly, this approach does not involve any loss of generality since in our high-level description based on the RWA, the scattering matrix is fully determined by the dimensionless BdG Hamiltonian $H$ (see Appendix F).
The importance of the rescaling in Eq. 2 is two-fold. On one hand, the weights $H_{ij}$ acquire a universal implementation-independent meaning as they are now directly related to the dimensionless physical parameters that determine the scattering behavior, i.e. the dimensionless detuning $(\omega_{L,i}-\omega_{i})/\kappa_{i}=-H_{ii}$, the cooperativities $C_{ij}=4|H_{ij}|^{2}$, and the gauge invariant phases accumulated on closed loops (synthetic field fluxes), $\Phi_{i,j,l,\ldots,k}=\arg(H_{ij}H_{jl}\ldots H_{ki})$. On the other hand, it implies that the algorithm discovers classes of solutions that all obey the desired target scattering behavior. When translating back to a concrete physical setup, this leaves the freedom to choose arbitrarily the decay rates $\kappa_{i}$ and, thus, fix the couplings accordingly (although for amplifiers some additional stability constraints might apply, see Appendix G).
For concreteness, we focus below on the important special case of phase-preserving devices. In this setting, the equations of motion for $(\hat{a}_{1},\ldots,\hat{a}_{N},\hat{a}_{1}^{\dagger},\ldots,\hat{a}_{N}^{\dagger})$ break down into two decoupled sets of equations which are related to each other via the particle-hole symmetry. As a result, the scattering matrix is an $N\times N$ matrix. For phase-preserving amplifiers, we enforce the block structure of the equations of motion by following the common engineering principle of dividing the modes (graph nodes) into two disjoint sets [39, 2, 31, 33, 37]. The modes within the same set are then only coupled via beamsplitter couplings, while the modes in different sets are only coupled via squeezing interactions. In this way, the annihilation operators for the modes in one set are only coupled to the creation operators of the modes in the other set (see Appendix F for more details). This constraint restricts the set of graphs considered in our optimisation. For example, a triangle with just one blue edge will be excluded. We do not indicate explicitly which mode belongs to the same group because this can be easily deduced from the colors of the edges (see Appendix F).
@@
Figure 2: Automated Discovery of Modular Scattering Setups. (a) Optimisation scheme behind AutoScatter. The discrete optimization sorts graphs into the libraries of valid (green book) and invalid graphs (red book). In every round, a new graph is suggested and, depending on the outcome of the continuous optimisation, sorted with all its extensions/subgraphs into the library of valid/invalid graphs. The continuous optimization minimizes the loss function to realize a target scattering behaviour (left block), here an isolator. The candidate graph (right block) defines the number of auxiliary modes and an allowed set of non-zero couplings rates and phases in the dimensionless Hamiltonian $H$. (b) Search space (schematic). The search starts from the smallest fully connected graph realising the target behavior (topmost graph). From there on, our method successively prunes edges, resulting in multiple irreducible solutions. (c) Comparison of the number of tested graphs for the directional coupler (see Fig. 4(b)), which uses five modes. The total number of possible graphs (black) for a certain number of free variables in $H$ is compared to the number of graphs which had actually to be tested by our exhaustive search. Our scheme starts with maximally connected graphs (left end) and successively goes to more restricted graphs (right end). By testing only around 13,000 graphs we are able to characterize all 1.9 million graphs in this example.
II Optimisation scheme
@@
To identify setups that implement the target scattering matrix, AutoScatter has to explore the discrete search space of all possible graphs and, given a graph, find an appropriate set of values for the underlying continuous parameters, i.e., the non-zero cooperativities and coupling phases. To this end, AutoScatter performs a two-step procedure: A discrete optimization routine suggests new graphs and an embedded continuous optimization algorithm looks for appropriate values of the parameters (see Fig. 2(a)). If the continuous optimization is successful, we label a graph as valid, i.e., this graph fulfills the desired scattering behavior.
@@
First, we discuss the discrete optimization, see Fig. 2(a). We aim to identify the simplest valid graphs as they represent the simplest experimental implementations of the target response. Often there are a multitude of valid graphs that differ in the number of auxiliary modes and the number and complexity of the required couplings. In a first step, AutoScatter tests only fully connected graphs for validity starting from the graph with zero auxiliary modes. It increments the number of auxiliary modes until it finds the minimum number of auxiliary modes required to achieve the target behavior. This is the smallest fully connected graph that satisfies the target behavior. From then on, AutoScatter prunes this fully connected graph and identifies “irreducible” graphs that can not be further simplified by setting any coupling rate or phase to zero. To identify all such graphs, we perform a complete enumeration.
@@
Due to the large discrete search space, a brute-force approach testing all graphs is infeasible in many cases of interest. We solve this problem by leveraging the knowledge gained from previously tested graphs, drastically reducing the required computational resources. We note that an alternative approach could be developed based purely on continuous optimization, introducing terms in a loss function that reward sparsity (zero couplings) as in [45, 46, 47] – however, this would not enable us to obtain a complete set of solutions.
@@
We take advantage of the fact that removing an edge from a graph is equivalent to setting a coupling rate to zero. Thus, if a graph is found to be invalid, that carries over to all graphs, which can be generated by removing some edges and/or setting coupling phases to zero. Likewise, all extensions of a valid graph are valid (see Fig. 2(b)). This observation allows us to efficiently construct exhaustive libraries of valid and invalid graphs by testing only a small fraction, see Fig. 2(c). In Section VII, we give an in-depth analysis of the search space reduction provided by AutoScatter. See also Appendix A for further details on the discrete optimization.
@@
We now turn to the continuous optimization. The scattering matrix $S$ is fully determined by the dimensionless Hamiltonian $H$ defined in Eq. 2, which contains the free coupling parameters. We look for setups implementing the desired scattering behavior by minimizing the squared deviation loss function ${\cal L}$ with respect to those couplings:
@@
We note that the target scattering matrix $S^{\mathrm{target}}$ can contain further free parameters that are not specified by the desired scattering properties. For example, we might impose that two scattering amplitudes are equal, but a suitable value needs to be discovered. In addition, we introduce the free gauge parameters $\gamma^{\rm in/out}_{i}$. This reflects the freedom of choosing independently for each port a reference quadrature and a reference position where the fields are measured. For more details, including how to fix the gauge in specific experimental settings, see Appendix J. Finally, we allow for extra constraints $f_{j}=0$ that can be used, e.g., to enforce minimum added noise as detailed in Appendix H. We minimize the loss ${\cal L}$ as a function of the free parameters in $H$, $S_{\mathrm{target}}$, and the phases $\gamma^{\rm in/out}_{i}$, see Appendix B for more details. A valid solution is obtained whenever the loss reaches zero.
Figure 3: Flowchart showing the full design pipeline of coupled-mode systems. The steps automatized by AutoScatter are highlighted in orange. The remaining steps, not covered by our method, are hardware-specific, see Section VIII.
The optimization described within this section is the main part of AutoScatter. In addition, AutoScatter provides an automated analysis of the complexity of the discovered solutions, as discussed in Section VIII. The broader design pipeline for a specific experimental platform including also steps not covered by AutoScatter is shown in the flowchart in Fig. 3.
In the following, we demonstrate through a series of illustrative examples how AutoScatter can design scattering setups, some of which have been overlooked by hand-crafted methods. We also show that our approach can automate several key aspects of scientific discovery. These include the generalization of design patterns, the finding of explicit analytical expressions for continuous classes of problems, and the discovery of asymptotic solutions.